OpenStudy (anonymous):

3.2 - 9x + 7.1 -3x help

OpenStudy (the_fizicx99):

3.2 - 9x + 7.1 - 3x 3.2 + 7.1 = 10.3 -9x -3x = -12x 10.3 - 12x

OpenStudy (anonymous):

i am helping my son with the same stuff. i found an algebra calculator that shows work at www.algebra.com. Or, you can google search an algebra calc that shows work. If you think you may know the answer from your multiple choice options, plug that number in the calc to see if it is right

OpenStudy (anonymous):

yes please

OpenStudy (the_fizicx99):

:3 or you can ask it here :)

OpenStudy (anonymous):

exactly!